f you have a circuit consisting of 4

4

 resistors of equal resistance in series and the total voltage across all of them is 17

17

 V

V

, what is the voltage across each of them?


1
Expert's answer
2022-06-09T09:59:11-0400

Given:

"R_1=...=R_4=R"

"V=17\:\rm V"


For the series connection

"V=V_1+V_2+V_3+V_4=4V_1"

"V_1=\frac{V}{4}=\frac{17}{4}=4.25\:\rm V"